OpenStudy (anonymous):

Find the sum of the infinite geometric series. -124/4 + 11/2 -1 + 2/11 -...

OpenStudy (anonymous):

r u sure its geometric series??

OpenStudy (anonymous):

thats is what is on my homewwork

OpenStudy (anonymous):

because according to me, first term should be -121/4

OpenStudy (anonymous):

your right

OpenStudy (anonymous):

so first term= -121/4 , common ratio= -2/11 hence sum= -121/4 /1+2/11 =-121/4/13/11 =-121/4 * 11/13

OpenStudy (anonymous):

so multiply that to get the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-1331/52 is what i get

OpenStudy (anonymous):

what ur book says?

OpenStudy (anonymous):

yep! thanks

OpenStudy (anonymous):

oh good .....keep up the good work